OpenStudy (anonymous):

5 - 3(2x + 4) = 25 After distributing, the question should read: 5 - 6x + 4 = 25 5 - 6x + 12 = 25 5 - 6x - 12 = 25 -4x - 8 = 25

OpenStudy (hoblos):

3(2x + 4) = 6x +12 so 5 - 3(2x + 4) = 25 5 - (6x+12) = 25 5 - 6x - 12 = 25